OpenStudy (pixiedust1):

Simplify 1*x-x/1 Please solve and show your work

OpenStudy (solomonzelman):

YOU MEANT SIMPLIFY. \[1 \times x = x\]so\[x-x/1\]\[x^2/x-1/x\]\[\frac{x^2-1}{x}\]

OpenStudy (pixiedust1):

So you're saying that the answer is 0?

OpenStudy (solomonzelman):

I am not saying that the answer is zero, I simplified it. But perhaps you meant, \[1 \times x-1/x=0\]

OpenStudy (pixiedust1):

Thank you. The options for my answers are... 0 1 X

OpenStudy (solomonzelman):

what is X, no solution?

OpenStudy (pixiedust1):

I think it means that it can only simplify down to x

OpenStudy (pixiedust1):

Are you still there?